Circular No. 6490

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The High-Z Supernova Search Team reports the discovery of seventeen

supernovae on CCD images taken at the Victor Blanco 4-m telescope at Cerro Tololo Interamerican Observatory (CTIO) by N. Suntzeff, A. Clocchiatti, and R. Schommer (CTIO); and B. Schmidt (Mount Stromlo and Siding Spring Observatories). Each supernova is within 2′′ of its host galaxy’s nucleus, except for SN 1996bg, which is 3′′.9 south and 0′′.3 west of its host galaxy, and all objects were observed at more than one epoch. The following are given for each supernova: the date of their discovery image, position, R magnitude at discovery, redshift, supernova type, and telescope (T) that provided the redshift and spectral-typing information.

SN  1996 UT  α_{2000}  δ_{2000}  R  z type T
1996at Sept. 15 1^{h}08^{m}17.^{s}63 −0^{o}58′25′′.5 20.5 0.09 Ic  A
1996au Sept. 15 3 12 47.79 +0 45 11.9 22.5 0.52 Ia  K
1996av Sept. 15 1 53 54.22 +0 15 30.6 22.4 0.51 Ia  K
1996aw Oct. 8 21 53 47.10 +0 02 47.3 21.7 0.21 Ia  M
1996ax Oct. 8 21 53 52.93 −0 02 36.4 23.3 0.23 ?  M/K
1996ay Oct. 8 0 09 57.41 +0 53 37.9 23.9 0.62 Ia  K
1996az Oct. 8 0 15 10.84 +1 06 34.6 24.0 0.84 ?  K
1996ba Oct. 8 1 08 39.87 −0 56 38.9 23.5 0.66 Ia  K
1996bb Oct. 8 2 57 06.06 −2 42 17.4 23.6 0.56 Ia?  K
1996bc Oct. 8 3 12 06.58 +0 38 17.1 22.5 0.46 Ia  K
1996bd Oct. 9 23 59 26.98 −0 56 16.6 21.6 0.28 Ia  M
1996be Oct. 9 2 12 02.83 +1 22 39.1 22.2 0.22 II  K
1996bf Oct. 9 2 12 15.39 +1 27 48.9 21.5 0.25 Ia  M
1996bg Oct. 9 0 52 52.38 −0 04 31.1 22.0 0.12 II M/K
1996bh Oct. 9 1 11 31.53 −0 57 38.1 22.8 0.43 Ia  K
1996bi Oct. 9 1 52 35.84 +0 02 43.3 23.0 0.60 Ia  K
1996bj Oct. 9 2 11 49.53 +1 24 51.0 22.6 0.57 Ia  K

The Anglo-Australian Telescope (A) observations were made by R. Stathakis (Anglo-Australian Observatory) on Oct. 9–10; The Keck Telescope (K) observations were made Oct. 11–12 by A. V. Filippenko, D. C. Leonard, and A. J. Barth (University of California, Berkeley); and the Multiple-Mirror Telescope (M) observations were made on Oct. 10–12 by R. Kirshner, Harvard-Smithsonian Center for Astrophysics. Finder charts may be requested by sending electronic mail to pgarnavich@cfa.harvard.edu.

1996 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green